OpenStudy (anonymous):

Write the point-slope form of an equation for a line that passes through the point with the given slope. (6, 3), slope = –1

OpenStudy (helder_edwin):

the equation would be (in general terms) \[ y-y_0=m(x-x_0) \] replace what you have and tell me what you get

OpenStudy (anonymous):

choices are a. y + 3 = –1(x – 6) c. y – 3 = –1(x + 6) b. y – 3 = (x – 6) d. y – 3 = –1(x – 6)

OpenStudy (helder_edwin):

ok but what do you get?

OpenStudy (anonymous):

y – 3 = –1(x + 6)

OpenStudy (helder_edwin):

you have m=-1 and \[ (x_0,y_0)=(6,3) \]

OpenStudy (helder_edwin):

you have a sign wrong